Measures of liquidity, Solvency and Profitability

The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall Inc. common stock was $ 71 on December 31, 20Y2.

Marshall Inc.
Comparative Retained Earnings Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Retained earnings, January 1 $ 2,461,325 $ 2,069,975
Net income 570,000 424,000
Total $ 2,966,025 $ 2,493,975
Dividends
On preferred stock $ 7,000 $ 7,000
On common stock 25,650 25,650
Total dividends $ 32,650 $ 32,650
Retained earnings, December 31 $ 2,998,675 $ 2,461,325


Marshall Inc.
Comparative Income Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Sales $ 3,919,370 $ 3,611,160
Cost of goods sold 1,312,540 1,207,540
Gross profit $ 2,606,830 $ 2,403,620
Selling expenses $ 935,880 $ 1,140,250
Administrative expenses 797,240 669,670
Total operating expenses 1,733,120 1,809,920
Income from operations $ 873,710 $ 593,700
Other income 45,990 37,900
$ 919,700 $ 631,600
Other expense (interest) 272,000 149,600
Income before income tax $ 647,700 $ 482,000
Income tax expense 77,700 58,000
Net income $ 570,000 $ 424,000


Marshall Inc.
Comparative Balance Sheet
December 31, 20Y2 and 20Y1
   Dec. 31, 20Y2    Dec. 31, 20Y1
Assets
Current assets
Cash $ 750,890 $ 607,950
Marketable securities 1,136,480 1,007,460
Accounts receivable (net) 686,200 642,400
Inventories 511,000 394,200
Prepaid expenses 142,059 121,590
Total current assets $ 3,226,629 $ 2,773,600
Long-term investments 897,589 13,273
Property, plant, and equipment (net) 4,420,000 3,978,000
Total assets $ 8,544,218 $ 6,764,873
Liabilities
Current liabilities $ 1,075,543 $ 1,363,548
Long-term liabilities
Mortgage note payable, 8 % $ 1,530,000 $ 0
Bonds payable, 8 % 1,870,000 1,870,000
Total long-term liabilities $ 3,400,000 $ 1,870,000
Total liabilities $ 4,475,543 $ 3,233,548
Stockholders' Equity
Preferred $ 0.70 stock, $ 50 par $ 500,000 $ 500,000
Common stock, $ 10 par 570,000 570,000
Retained earnings 2,998,675 2,461,325
Total stockholders' equity $ 4,068,675 $ 3,531,325
Total liabilities and stockholders' equity $ 8,544,218 $ 6,764,873

Required:

Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year.

Return on stockholders’ equity %
Return on common stockholders’ equity %
Earnings per share on common stock $
Price-earnings ratio
Dividends per share of common stock $
Dividend yield
Answer #1
1. Return on Stockholder's quity: Net income after tax/ Average stockholders equity
Average Stockholder equity (3531325+4068675)/2= $3800,000
Net income after tax: $ 570000
Return on Stockholders' equity: 570000/3800,000 *100 = 15%
2. Return on Common Stocllholder' equity: Net income after preferred div / Averagee common Stoclholder sequity
Average common Stocholder's equity : 3800,000 - 500,000 = 3300,000
Net income after preferred dividend: 570000-7000 = 563000
Return on Common Sstockholder's equity: 563000 /3300,000*100= 17.06%
3.Earnings per share: Net income after dividend / Number of shares
563000 /57000 = $9.88 per share
4. Price earning ratio: market prikce per share/ Earning per share
71 / 9.88 = 7.19
5. Dividend per share of common stock: Total dividend/ Number of common Shares
25650 / 57000 = $ 0.45 per share
6. Dividend yield ratio: Dividend per share/ market price per share*100
0.45/ 71 *100 = 0.63%